# MARCH: Evaluating the Intersection of Ambiguity Interpretation and Multi-hop Inference

### v2

@@ 6 Related Work
Multi-hop QA requires reasoning across multiple documents (he2024mintqa; zhu2024fanoutqa; tangmultihop). HotpotQA (yang2018hotpotqa) targets the retrieval of articles and sentence-level facts, while 2WikiMultihopQA (ho-etal-2020-constructing) enhances explainability by providing structured evidence and reasoning paths. MuSiQue (trivedi2022musique) mitigates shallow shortcuts found in prior datasets by enforcing connected reasoning through dependent single-hop questions. Unlike prior ambiguity benchmarks (e.g., AmbigQA, ASQA, CondAmbigQA (li2025condambigqa)), which target single-hop questions, MARCH targets the intersection of ambiguity and multi-step inference. As summarized in Table 5, MARCH evaluates the entire pipeline of multi-hop ambiguity detection, clarification, and answer generation.
@@ 7 Conclusion
CLARION demonstrates strong performance but introduces additional system complexity due to its multi-agent structure and planning-acting cycle. Although this complexity enables richer ambiguity resolution, future work could explore lighter-weight or more efficient designs without sacrificing effectiveness, making deployment and integration even more practical.
@@ Acknowledgments
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University)]. This research was supported by the Chung-Ang University Graduate Research Scholarship in 2025.
@@ Appendix D Latency Analysis
Appendix E Results on Other Related Benchmarks
E.1 Results for ASQA Benchmark
E.2 Results for MuSiQue Benchmark
Method Qwen3-235B Gemini-2.5-Flash DeepSeek-Chat-V3.1
EM F1 EM F1 EM F1
Naive RAG 0.040 0.558 0.102 0.520 0.044 0.478
DIVA 0.142 0.576 0.106 0.520 0.106 0.518
ReAct Only 0.140 0.568 0.132 0.551 0.132 0.509
Clarion 0.206 0.594 0.146 0.560 0.160 0.514
Table 12: Comparison of EM and F1 scores across different models and methods.
Appendix F LLM-as-a-Judge
F.1 Validating the LLM Judge.
@@ Appendix F LLM-as-a-Judge
F.2 Detailed Score for LLM-as-a-Judge
@@
Appendix G Case Study
G.1 Failure Cases of CLARION
@@
G.2 Case Study: A 3-hop Query with Three Interpretations
@@
Appendix H Implementation Details
@@
All LLM calls (for both detection and generation) were run with temperature set to 0.0 and a maximum token limit of 512. All prompts and task templates are described in detail in Section J.
@@
Appendix I Detection cue for Constraint ambiguity
@@
Appendix J Prompt Templates

### v4

@@ 6 Related Work
Multi-hop QA requires reasoning across multiple documents (He et al., 2024; Zhu et al., 2024; Tang and Yang, 2024). HotpotQA (Yang et al., 2018) targets the retrieval of articles and sentence-level facts, while 2WikiMultihopQA (Ho et al., 2020a) enhances explainability by providing structured evidence and reasoning paths. MuSiQue (Trivedi et al., 2022) mitigates shallow shortcuts found in prior datasets by enforcing connected reasoning through dependent single-hop questions. Unlike prior ambiguity benchmarks (e.g., AmbigQA, ASQA, CondAmbigQA (Li et al., 2025)), which target single-hop questions, MARCH targets the intersection of ambiguity and multi-step inference. As summarized in Table 5, MARCH evaluates the entire pipeline of multi-hop ambiguity detection, clarification, and answer generation.
@@ 7 Conclusion
CLARION demonstrates strong performance but introduces additional system complexity due to its multi-agent structure and planning-acting cycle. Although this complexity enables richer ambiguity resolution, future work could explore lighter-weight or more efficient designs—for instance, a hybrid routing strategy that reserves CLARION for complex, ambiguous queries while forwarding simpler ones to a standard single-step pipeline—without sacrificing effectiveness, making deployment and integration even more practical.
@@ Acknowledgments
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)] and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2026-25494299). This research was supported by the Chung-Ang University Graduate Research Scholarship in 2025.
@@ C.2 Human Evaluation Protocols
Method Qwen3-235B Gemini-2.5-Flash DeepSeek-Chat-V3.1
EM F1 EM F1 EM F1
NaiveRAG 0.040 0.558 0.102 0.520 0.044 0.478
DIVA 0.142 0.576 0.106 0.520 0.106 0.518
ReAct Only 0.140 0.568 0.132 0.551 0.132 0.509
CLARION (ours) 0.206 0.594 0.146 0.560 0.160 0.514
Table 13: MuSiQue results. Comparison of EM and F1 scores across different models and methods.
@@ Appendix D Latency Analysis
Appendix E LLM-as-a-Judge
E.1 Validating the LLM Judge.
@@ Appendix E LLM-as-a-Judge
Appendix F Results on Other Related Benchmarks
F.1 Results for ASQA Benchmark
F.2 Results for MuSiQue Benchmark
F.3 Detailed Score for LLM-as-a-Judge
@@
Appendix G ReAct Iteration Limit Ablation
We set the maximum number of ReAct iterations to five based on an empirical ablation varying the limit across $\{1,3,5,7\}$ on Qwen3-235b. As shown in Table 15, performance improves steadily as the iteration limit increases from 1 to 5. Beyond 5 iterations, however, gains saturate: increasing the limit to 7 yields only negligible improvement (++0.10 Avg) while incurring unnecessary computational overhead and latency. We therefore adopt 5 as the optimal trade-off between performance and efficiency.
Iter. STR-EM Disambig-F1 Avg LLM-Judge
1 28.86 22.45 25.65 2.850
3 36.82 26.91 31.86 3.315
5 38.73 28.38 33.56 3.474
7 38.81 28.52 33.66 3.461
Table 15: Impact of ReAct iteration limit on CLARION (Qwen3-235b). Performance saturates beyond 5 iterations with negligible gains. Bold = best; underline = second best.
Appendix H Case Study
H.1 Failure Cases of CLARION
@@
H.2 Case Study: A 3-hop Query with Three Interpretations
@@
Appendix I Implementation Details
@@
All LLM calls (for both detection and generation) were run with temperature set to 0.0 and a maximum token limit of 512. All prompts and task templates are described in detail in Section K.
@@
Appendix J Detection cue for Constraint ambiguity
@@
Appendix K Prompt Templates